I am currently 60 years old. My grand daughter is 12. At some time I will be three times her age. Write an equation which models how old each of us will be.
Found 2 solutions by jorel1380, solver91311:
Answer by jorel1380(3719)

60+n=3(12+n)
60+n=36+3n
2n=24
n=12
In 12 years you will be 3 times your granddaughter's age (72 and 24)..

Answer by solver91311(24713)



Let represent the number of years between now and the time the grandfather is three times as old as the granddaughter.

Then



In plain English: Three times the age the grandaughter will be years from now is the same as the age the grandfather will be years from now.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism